A circle and a hyperbola can have a maximum of how many intersection?
1
Expert's answer
2019-12-09T11:00:27-0500

We need to find the maximum number of intersections of a circle and a hyperbola.


The Equation of a circle is

"x^2 + y^2 = r^2 ...\u2026\u2026......\u2026..(1)"

The equation of a hyperbola is

"\frac {x^2}{a^2} - \frac {y^2}{b^2} =1 ………...……….(2)"

The intersection points can be obtained as the solution of both the above equations.


From equation (1),

"y^2 = r^2 - x^2"

Plug this in equation (2),

"\frac {x^2}{a^2} - \frac {r^2 - x^2}{b^2} =1"

"b^2 x^2 - a^2 r^2 + a^2 x^2 = a^2 b^2"

"x^2 ( a^2 + b^2 ) = a^2 b^2 + a^2 r^2"


"x^2 = \frac {a^2 (b^2 + r^2)}{a^2 + b^2}"


"x = ± \sqrt {\frac {a^2 (b^2 +r^2)} {a^2 + b^2}}"

Plug the "x^2 = \frac {a^2 (b^2 + r^2)}{a^2 + b^2}" in "y^2 = r^2 - x^2"



"y^2 = r^2 - \frac {a^2 (b^2 + r^2) }{a^2 + b^2 }"

"y^2 = \frac {a^2 r^2 + b^2 r^2 - a^2 b^2 - a^2 r^2}{a^2 + b^2}"

"y^2 = \frac {b^2 (r^2 - a^2)}{a^2 + b^2}"

"y = ± \sqrt {\frac {b^2 (r^2 - a^2 )} {a^2 + b^2}}"

So, the intersecting points are pairs (x,y) taking the sign into account just like (+, + ), (+, -), (-, + ), ( -, -).


The maximum number of intersections is 4.


Answer: the maximum number of intersecting points is 4.